Shift report 20030311 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Mar 11 2003 - 00:00:08 EST

  • Next message: Apache: "Shift report 20030311 00:00-08:00"
    ------------------------------------------------------------
    Date: Tue Mar 11 00:00:08 2003, Shift: 16:00-24:00
    Supervisor: FR, with: SL
    Log book#14 pages 55 - 57
    
    Shift Summary:
    ------------------------------------------------------------
    
    1) After a couple of unsuccessful ramps, MCR could store
       the beam from 20:20 to 21:10. They dumped for a RHIC
       cryo ring access. Next store started around 23:40
       with initial rates: ZDC~11k and TOFW~18k.
    
    2) Runs 8304-8306 are calibration runs.
       Run 8307 is a data run for MRS @90deg - 1050A and
       FS @4deg - 1/1A. It has ~3k trigger 2 events.
       Run 8308 started at the end of the shift (same setting
       as Run 8307).
    
    3) The inelastic trigger was retimed by Flemming (see
       Log Book 14, pg 55-56 for details)
       No detector problems.
    
    ------------------------------------------------------------
    


    This archive was generated by hypermail 2.1.5 : Tue Mar 11 2003 - 00:00:56 EST